OpenStudy (anonymous):

The graph of the following system of equations is: -2x + y = 3 4x + 2y = 2 Answers- Overlapping lines Parallel lines Intersecting lines

OpenStudy (anonymous):

Is it intersecting lines? I plugged them into Desmos but I'm unsure

OpenStudy (victoriasushchik):

As all equations are written in the same form ax+by=c just find a/b (coefficient of x/coefficient of y) in all cases (or pairs) wherever a/b is identical the pair of lines are parallel for eg. in case (a) a/b for line x-2y = 3 is 1/-2 = -2 and for x-2y =1 a/b is again 1/-2 = -2 hence the pair indicated are parallel in the above cases except for the pair of lines indicated in (d) all the other pairs are parallel (question (c) is a typo where z should be x) This is the fastest way of finding out you dont need to rearrange the equation to find slopes. visual finding of a/b is possible theory is slope of line ax +by +c = 0 is (-a/b) . hence finding a/b and comparing is enough as all lines are in same form. lines with equal slopes are parallel